nitilly at rest experiences an acceleration of 2.00 m/se for 5.10 s then travels at that constant velocity for another 9 50s. What is the magnitude of the objects average velocity over the 14.6 s interval? m/s
Answer #1

First distance,

d1 = 0.5 x 2 x 5.10^2 = 26.01 m

Speed after 5.10 sec,

V = 2 x 5.1 = 10.2 m/s

d2 = 10.2 x 9.50 = 96.9 m

Average speed, v_avg = (96.9 + 26.01)/14.6 = 8.42 m/s

Comment in case any doubt please rate my answer .....